Jesper Co begins its next year of operations with a Beginning Supplies balance of 1,400. The end of period count of inventory revealed an Ending Supply balance of 4,500. If COGS for the same period was 36,700, Purchases must have been what?

36,700

33,600

35,300

39,800

On April 1 the ABC Inc. paid $20,400 to Acme Realty for 12 months Rent beginning May 1. The company did the correct recording on April 1. If financial statements are prepared on December 31st, the rent expense for the year is:

15,300

13,600

11,900

20,400
Answer #1

Cost of Goods Sold = Beginning inventory + Purchases - Ending inventory

36,700 = 1,400 + Purchases - 4,500

Purchases = 36,700 - 1,400 + 4,500

= 39,800

Rent expense for the year

= (20,400/12 months) * 8 months (May to Dec)

= 1,700 * 8

= 13,600